Hi I have a couple questions but I think I got the right answer but can you show step by step

e the eld ls trlanging The plane of a square loop of wire with edge length of 8.0 cm is perpendicular to a 0.017 T magnetic field. What is the average EMF in the loop if it is folded on the diagonal (so thatthe area of the loop goes to Oj ifit takes 0.13 sto make the fold? A flip coil of 50 mm radius and 50 turns is oriented perpendicular to a magnetic field. The coil is quickly flipped through 180 degrees about its diameter. It takes 0.070 s to flip the coil. The average EMF in the coil is measured to be 2.0 V. What is the strength of the magnetic field in the region of the coi? DOLL Shen
Answer #1

delta A = -0.08 * 0.08 = -0.0064 m2

emf = - delta (B A) / delta T

emf = - B * delta A / delta t

= (-0.017 * -0.0064) / (0.13)

= 8.37 * 10-4 V